OpenStudy (anonymous):

Find the perimeter of an equilateral triangle with side length of 12 inches.

OpenStudy (anonymous):

The rectangle will have a height of a and a width of b. Call the two line segments to the left and right of the base of the rectangle s. They are equal, you can prove this quite straightforwardly. So, 2s + b = 12. Also s = a /tan60 substituting for s, 2a /tan60 + b = 12 2a = (12 - b)tan60 a = (12-b) tan60/2 the area of the rectangle is ab or ab = (12 - b) * b * tan60/2 Clearly, the max value of this doesn't require the constants at the end, so we need to find the max value of 12b - b squared. You can do this in a variety of ways, but the answer is 6. Solving for a, it's also 6. So the max rectangle is a square of side 6

OpenStudy (anonymous):

awwhhh thank you! You just helped me a lot..(:

OpenStudy (anonymous):

Glad I did, if ya ever need me you got me (:

OpenStudy (anonymous):

awh thankss i think i like you already! haha

OpenStudy (anonymous):

I think I like ya too haha

OpenStudy (anonymous):

well dont i feel special!(:

OpenStudy (anonymous):

Nahh don't cuz I'm already feeling special ^.^

OpenStudy (anonymous):

haha you got me blushin n i dont even know you! Lol

OpenStudy (anonymous):

Now I'm blushing man, I blame it on you though!

OpenStudy (anonymous):

Lol message meee..

OpenStudy (anonymous):

In a flash